OpenStudy (anonymous):

What form of a linear equation is most useful when using the elimination method for solving a system of linear equations? Why?

OpenStudy (jdoe0001):

have you used the elimination method with a system of equations yet?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

so when solving a system of equations, is easier if one puts all the variables on one side, usually the left, and the constant value on the right

OpenStudy (jdoe0001):

say 3x + 25 = 2y y + 3 = 5x would be better managed if 3x -2y = -25 5x +y = -3

OpenStudy (anonymous):

yes, but why is it easier that way?

OpenStudy (jdoe0001):

http://www.mathsisfun.com/algebra/standard-form.html <--- you'd notice that'd be the "standard form of a linear equation"

OpenStudy (jdoe0001):

well.... when you've done it.... do you think it was simple? why?

OpenStudy (jdoe0001):

it's easier, since you can process vertically like you'd on any sum procedure